OpenStudy (anonymous):

There are 6 cans per box. Find the number of cans in 11 boxes. (Show Work) + Write the unit rate as a ratio. Then find an equivalent ratio.

jimthompson5910 (jim_thompson5910):

1 box = 6 cans 11*1 box = 11*6 cans 11 boxes = 66 cans So there are 66 cans in 11 boxes. An equivalent ratio to 1 box: 6 cans is 11 boxes: 66 cans